Shift report 20000730 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Jul 30 2000 - 16:15:52 EDT

  • Next message: http account: "Shift report 20000730 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Jul 30 16:15:51 2000, Shift: 08:00-16:00
    Supervisor: oh, with: +hi
    Log book#3 pages 115 - 118
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam till 1145 and again from 1515, b=500, y=120
    2. Runs with gating MTPC1=250+-80 V
       and with no gating on T1. D1 @ 520 A all the time.
       There appears to be a shift of -5 to 8 ns in the timing,
       that on might believe the intersection has moved from
       5 to 8 feet. Phobos does see such a move of IR. Effect
       is unexplained.
    3. TPC-1, MTPC-1 and MTPC-2 have nice tracks, TPC-2 none.
       T3 was reset to voltage, probably a trip.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Jul 30 2000 - 16:19:04 EDT